Follow this link to skip to the main content

Cassini Significant Events -- 05/18/06 - 05/24/06

Cassini Significant Events -- 05/18/06 - 05/24/06

May. 26, 2006


[ - ]   Text   [ + ]
Cassini Significant Events -- 05/18/06 - 05/24/06
May 26, 2006
(Source: Cassini Project)

The most recent spacecraft telemetry was acquired on Wednesday, May 24, from the Goldstone tracking complex. The Cassini spacecraft is in an excellent state of health and all subsystems are operating normally.

Friday, May 19 (DOY 139):

Per commands in the background sequence, the Main Engine (ME) cover was closed today as a preventive measure for dust avoidance near Dione's orbit. It will be reopened on Monday. The last time the cover was moved was about seven months ago. It was closed on October 9, 2005, for a ring plane crossing, and then reopened on October 12. It will not be moved again till late June of 2006.

Saturday, May 20 (DOY 140):

The big news today is the Titan 14 (T14) flyby. Cassini passed by the satellite at an altitude of 1879 kilometers. 

During this flyby, the Radio Science Subsystem (RSS) took advantage of an ingress/egress occultation of Earth by Titan to provide a high-spatial-resolution electron number density profile of Titan's ionosphere, temperature, pressure, and absorption profiles of the neutral atmosphere, as well as information about the small-scale structure of the atmosphere such as gravity waves, turbulence, etc. 

The Madrid DSN complex covered both the ingress and egress observations. The 70-m DSS-63 covered S- and X-bands. The 34-m DSS-55 covered Ka-band and provided backup X-band. RCP and LCP polarizations were recorded at the Earth receiving stations.

A "quick-look" analysis of the signal intensity profiles computed from the 1 KHz bandwidth recorded data indicate high quality S-, X-, and Ka-band ionospheric and atmospheric occultation data were acquired on both the ingress and egress sides. Signal frequency data was also collected and appears to be of equally good quality.

After a "quick-look" analysis of the recorded 16 KHz bandwidth T14 data analyzed last night, the RSS Team is happy to report the successful detection of the first ever bistatic echo received from the surface of Titan. It is a whisper, as if Titan is trying to hold back on one more of its secrets, but is no doubt real, and is beautiful! The detection is, of course, the first step towards the goal of finding out what the echo tells us about the surface region probed.

The similarities between the occultation observations at T12 and T14 indicate very little pointing errors, although a quantitative assessment is yet to be completed. The limb-track maneuver appears to have executed as planned. T12 and T14 are the only two RSS Titan occultations during 2006. Titan 27 and 31 will follow in March and May of 2007. The excellent data from T12 and T14 will keep the RSS team busy until T27!

The Composite Infrared Spectrometer (CIRS) performed high spectral resolution studies of Titan's limb in the far-infrared regime to search for new species and to map the vertical distribution of CO, CH4, HCN and H20. CIRS also continued existing campaigns of global temperature composition mapping, extending spatial and temporal coverage. 

The Ultraviolet Imaging Spectrograph (UVIS) probed haze optical properties, atmospheric nitrogen emissions and hydrocarbon absorption, and absorption by methane of the Lyman-alpha interplanetary background, which will ultimately help us understand the distribution of methane in the thermosphere of Titan.

The Cosmic Dust Analyzer (CDA) performed two eccentricity scans. These observations allow separation of orbital elements to the greatest degree possible by having the spacecraft rolling with a rotation axis which is parallel to the ring plane so that dust particle inclinations can be determined.

Sunday, May 21 (DOY 141):

On Sunday and Monday, UVIS performed some Saturn occultation observations using the stars Beta Orionis, Epsilon Orionis, and Zeta Orionis. These observations will yield the temperature of Saturn's high atmosphere and vertical profiles of H, H2, and hydrocarbons. The Visual and Infrared Mapping Spectrometer (VIMS) performed a couple of Saturn thermal cylindrical mappings as well.

At the end of the day today, there were some exhausted but very happy Cassini folks here at the lab. This weekend, JPL held its annual Open House. Approximately 30,000 people visited the lab during the two-day event. Cassini had a large exhibit on the JPL mall. The exhibit featured images from the Saturn tour, a display of Synthetic Aperture RADAR images, two plasma screen monitors showing Cassini video clips, a full-scale model of the Huygens Probe, and a half-scale model of the Cassini orbiter - including half scale people! The DVD "Ring World" was showing in a nearby conference room. Cassini image data and "sounds" were also featured in the DSN and 'Sounds of the Solar System' booths. A Girl Scout troupe used the "Moon Walk, Saturn Style" on their scavenger hunt to answer questions, work on badges and walk Saturn's moons from Pan to Phoebe. Feedback from the public was extremely positive. Nearly 40 members of the flight team spent all or part of their weekend visiting with guests at our locations.

Monday, May 22 (DOY 142):

A non-targeted flyby of the Saturnian satellite Polydeuces occurred today at an altitude of 64057 kilometers.

The ME cover opened today per plan after the T14 flyby, and after Cassini had cleared any possible dust hazards.

Tuesday, May 23 (DOY 143):

A delivery coordination meeting was held today for Inertial Vector Propagator (IVP) version 12.0, and Kinematic Prediction Tool (KPT) version 12.0. The software was approved and installation plans and instructions were reviewed.

The Spacecraft Team uplinked an RWA bias today in place of Orbit Trim Maneuver #62.

Wednesday, May 24 (DOY 144):

The sequence leads for S21 requested a loading test for the transition from S20 part 2 to the S21 background sequence. That test completed successfully today in the Integrated Test Laboratory. There should be no problems with the start of execution of S21.

It has been determined that the magnitude of Orbit Trim Maneuver (OTM) #62 would only be 0.8 mm/sec, and as a result, OTM 62 has been cancelled.

A delivery coordination meeting was held today for the tool Electronic Command Request Form version 1.5. The software was approved and installation plans and instructions were reviewed.